OPTICAL SPECTROSCOPY OF NEGATIVE IONS BASED ON LASER ASSISTED ELECTRON ATTACH-MENT

E. B. GORDON, and \underline{O. S. RZHEVSKY}, Institute for Energy Problems of Chemical Physics (branch), 142432 Chernogolovka, Moscow Region, Russian Federation.

Along with known processes of atoms and molecules photoassociation, another process of photoassociative electron attachment, A + e + hν → A^{−∗∗} → A + e , must exist for any particle, including those without positive electron affinity; its rate constant providing laser light flux of I ∼ 10^{8} W/cm^{2} is estimated by us to amount to 10^{−13}cm^{3}s^{−1}. The corresponding spectrum lies at energies of 10..15 eV and contains quantum state and structure information scarcely attainable theoretically.

The feasibility of developing universal technique for obtaining the short-lived negative ion spectra is discussed. It is based on the following considerations:

− since the energy of an ion state is a sum of those of a photon and an impact electron, the KrF excimer laser (hν = 5 eV) can be used, whereas the tuning is implemented by the electron beam energy variation in the range of ε = 5..10 eV; − since the energy of a secondary electron arising after a negative ion decay exceeds that of an impact electron by the value of hν, secondary electrons could be separated easily by the methods of electron optics; − moreover, in a number of cases the spatial distribution of secondary electron velocities correlates with the laser light polarization; − impact electrons having the energy of several eV can penetrate into the high electron density regions of target atoms and molecules , so that the probability of photoassociative electron attachment seems to be considerable even in the case of species with negative electron affinity.